OpenStudy (anonymous):

Company A rents copiers for a monthly charge of $300 plus 12 cents per copy. Company B rents copiers for a monthly charge of $600 plus 6 cents per copy. What is the number of copies above which Company A's charges are the higher of the two?

OpenStudy (anonymous):

let the number of copies above which company A's charge is higher be x company A's cost for x copies = 300 + 0.12 x company B's cost for x copies = 600 + 0.06 x there's a point at which company A's cost is equal to company B's cost, after which company A's cost becomes more than company B's cost. let's find that point 300 + 0.12x = 600 + 0.06x 0.12x - 0.06x = 600 - 300 0.06x = 300 x = 300/0.06 = 30000/6 = 5000 so, above 5000 copies, company A's charge is higher